OpenStudy (mrs.ambrose614):

Question 15 (Multiple Choice Worth 1 points) (03.05 LC) The graph of a function f(x) is shown below: plot ordered pairs negative 1, negative 3 and 0, 1 and 1, 2.2 and 2, 3 What is the domain of f(x)? integers from −1 < x < 2 integers from −3 < y ≤ 3 integers from −3 ≤ y < 3 integers from −1 ≤ x ≤ 2 Question 16 (Multiple Choice Worth 1 points) (03.03 MC) Choose the equation that represents the graph below: Graph of a line passing through points negative 2 comma 0 and 0 comma negative 4 y = −2x − 4 y = 2x + 4 y = −2x + 4 y = 2x − 4 Question 17 (Ye

OpenStudy (will.h):

The domain is the x values in any point. Therefor it would be −1 ≤ x ≤ 2 The other question: (-2,0)(0,-4) the y intercept is -4 because that's when x = 0 (as we previously discussed) and all we need to do is find the slope using : y2-y1/x2-x1 Slope = -4/2 Slope = -2 so -2x - 4